OpenStudy (anonymous):

A photograph with the width of 5inches and a length of 7inches will be enlarged so that the length is 56 inches.What will the width equal? A)35inches B)40inches C)92inches D)180inches

OpenStudy (anonymous):

You have two similar rectangles which means you can set the sides in a ratio and make them equal For Example Rectangle ABCD and Rectangle EFGH \[\frac{AB}{EF} = \frac{CD}{GH}\]

OpenStudy (anonymous):

5/7= x/56 just cross multiply and you will get x

OpenStudy (anonymous):

Your question is a ratio problem